# Hybrid Beamforming Design for Near-Field ISAC with Modular XL-MIMO

### v1

@@ III-C Problem Formulation
In this paper, we jointly design the analog beamformer $\mathbf{W}_{\text{RF}}$ and digital beamformer $\mathbf{W}_{\text{BB}}$ in order to maximize the achievable communication spectral efficiency while satisfying the constraints imposed by the transmit power budget and sensing SCNR. Based on (14) and (20), the optimization problem can be given by
$\max\limits_{{{\mathbf{W}}_{\text{BB}}},{{\mathbf{W}}_{\text{RF}}}}$ $\ \ \ C$ (22a)
@@
$\ \ {{\gamma}_{s}}\geq{{\Gamma}_{s}},$ (22d)
where ${{\mathcal{A}}_{F}}$ is the set of block matrices, where each block is an $M\times M_{\text{RF}}$ dimension matrix with constant-magnitude entries. However, due to the specific block-diagonal structure of the analog beamformer in ${\mathcal{A}}_{F}$, this type of joint analog-digital optimization problem is often intractable, leading to the attainment of suboptimal solutions rather than a global optimal solution[15].
@@ IV-A The Optimal Waveform Covariance Matrix
To provide an optimal solutions to the problem (23), we propose to exploit the structure of the communication channel and sensing array response vectors based on HSPM, respectively. Namely, we leverage the following observations:
IV-A1 Structure of the communication channel based on HSPM
@@
IV-A2 Structure of Sensing Array Response Vectors based on HSPM
For a given sensing object $q$, the subarray response vectors $\mathbf{a}_{tq}^{1}(\mathbf{l}_{q}),\ldots,\mathbf{a}_{tq}^{K}(\mathbf{l}_{q})$ are linearly independent because different subarrays observe different AoD of $\mathbf{l}_{q}$. Therefore, each sensing array response vector ${{\mathbf{g}}_{t}}\left({{\mathbf{l}}_{q}}\right)$ can be expressed as a fixed linear combination of $K$ linearly independent vectors as
$\mathbf{\bar{a}}_{tq}^{k}{{\left(\phi_{tq}^{k}\right)}^{T}}\triangleq\left[\underbrace{{0},\ldots,{0}}_{k-1},\mathbf{a}_{tq}^{k}{{\left(\phi_{tq}^{k}\right)}^{T}},\underbrace{{0},\ldots,{0}}_{K-k}\right],\quad\forall k.$ (25)
@@ IV-A2 Structure of Sensing Array Response Vectors based on HSPM
${{\mathbf{g}}_{t}}\left({{\mathbf{l}}_{q}}\right)={{\mathbf{\bar{A}}}_{tq}}{\bm{\nu}}_{tq},\ \forall q,$ (26)
@@ IV-A3 Joint Representation of Communication Channel and Sensing Array Response Vectors
${{\mathbf{\tilde{U}}}_{k}}\!$ $=\!\left[\mathbf{\bar{a}}_{t1}^{k}\!\left(\phi_{t1}^{k}\right)\!,\ldots,\mathbf{\bar{a}}_{tQ}^{k}\!\left(\!\phi_{tQ}^{k}\!\right)\!,\mathbf{\tilde{a}}_{t1}^{k}\!\left(\theta_{t1}^{k}\right)\!,\ldots,\mathbf{\tilde{a}}_{t{{N}_{p}}}^{k}\!\left(\!\theta_{t{{N}_{p}}}^{k}\!\right)\!\right],$ (35a)
${{\mathbf{A}}_{kk}}\!$ $=\!\left[\mathbf{a}_{t1}^{k}\!\left(\phi_{t1}^{1}\right)\!,\ldots,\mathbf{a}_{tQ}^{k}\!\left(\phi_{tQ}^{1}\right)\!,\mathbf{a}_{t1}^{k}\!\left(\theta_{t1}^{k}\right)\!,\ldots,\mathbf{a}_{t{{N}_{p}}}^{k}\!\left(\!\theta_{t{{N}_{p}}}^{k}\!\right)\!\right].$ (35b)
@@ IV-B Equivalent Low-dimensional Optimization Problem
To fully utilize the multiplexing gain, we have $N_{s}=\mathrm{rank}(\mathbf{H}_{c})=KN_{p}$. In addition, in the mmWave/sub-THz band, where reflection and scattering losses are significant, the contributions of high-order reflection and scattering paths can be neglected, leading to a small value of $N_{p}$. Therefore, in scenarios where the number of subarrays and the number of interferences are limited, we can consider $N_{p}+Q$ to be much smaller than $M$. Consequently, we can use $M_{\text{RF}}=N_{p}+Q$ RF chains at each subarray, and the total number of RF chains is $N_{\text{RF}}=K(N_{p}+Q)$.
@@
${{\gamma}_{s}}=\frac{\alpha_{1}^{2}\text{tr}\left({{\mathbf{W}}_{\text{BB}}}\mathbf{W}_{\text{BB}}^{H}{{\mathbf{\Phi}}_{1}}\right)}{\sum\nolimits_{q=2}^{Q}{\alpha_{q}^{2}\text{tr}\left({{\mathbf{W}}_{\text{BB}}}\mathbf{W}_{\text{BB}}^{H}{{\mathbf{\Phi}}_{q}}\right)}+\sigma_{s}^{2}\mathbf{w}{{\mathbf{w}}^{H}}},$ (38)
where ${{\mathbf{\Phi}}_{q}}\triangleq{{\mathbf{\tilde{U}}}^{H}}{{\mathbf{g}}_{tq}}\mathbf{g}_{rq}^{H}\mathbf{w}{{\mathbf{w}}^{H}}{{\mathbf{g}}_{rq}}\mathbf{g}_{tq}^{H}\mathbf{\tilde{U}}$, $\forall q\in\left\{1,2,\ldots,Q\right\}$. Moreover, due to the block diagonal structure of $\mathbf{{W}}_{\text{RF}}$, each non-zero element of $\mathbf{W}_{\text{RF}}$ is multiplied with the corresponding row in $\mathbf{{W}}_{\text{BB}}$. Hence, the constraint (22b) can be simplified as
@@
[…]
$\begin{split}{}&\ \ \alpha_{1}^{2}\text{tr}\left(\mathbf{W}_{\text{BB}}^{H}{{\mathbf{\Phi}}_{1}}{{\mathbf{W}}_{\text{BB}}}\right)\text{-}{{\Gamma}_{s}}\sum\limits_{q=2}^{Q}{\alpha_{q}^{2}\text{tr}\left(\mathbf{W}_{\text{BB}}^{H}{{\mathbf{\Phi}}_{q}}{{\mathbf{W}}_{\text{BB}}}\right)}\geq{{\Gamma}_{0}},\\[-8.53581pt] \end{split}$ (40c)
where $\Gamma_{0}\triangleq{{\Gamma}_{s}}\sigma_{s}^{2}\mathbf{w}{{\mathbf{w}}^{H}}$.
Althogh the digital beamformer optimization problem (40), has a reduced dimension, it remains non-convex due to the rank constraint imposed by the limited number of data streams. To tackle this non-convex problem, we propose two distinct algorithms in the following subsections: a manifold optimization method that directly optimizes the digital beamformer on the rank-constrained space and an SDR-based method that obtains a near-optimal solution. By investigating these two optimization strategies, we provide a comprehensive framework for solving the rank-constrained digital beamformer design problem.
@@ IV-C Joint Optimization on Riemannian Manifold and Euclidean Space
In this subsection, we first analyze the structure of the optimal solution to problem (40) without relaxing the rank constraint, and obtain the semi-closed-form solution. Based on this, we then develop a Riemannian joint gradient descent algorithm.
Let $\mathbf{B}\triangleq{{\mathbf{\tilde{U}}}^{H}}\mathbf{H}_{c}^{H}{{\mathbf{H}}_{c}}\mathbf{\tilde{U}}$, and it follows that $\mathbf{B}$ is a Hermitian matrix with $\mathrm{rank}(\mathbf{B})=\min\{N_{\text{RF}},KN_{p}\}=N_{s}$. Performing eigendecomposition on $\mathbf{B}$ and retaining only the non-zero eigenvalues and their corresponding eigenvectors, we obtain $\mathbf{B}={{\mathbf{U}}_{B}}{{\mathbf{\Sigma}}_{B}}\mathbf{U}_{B}^{H}$, where ${{\mathbf{U}}_{B}}\in\mathbb{C}^{N_{\text{RF}}\times N_{s}}$, and $\mathbf{\Sigma}_{B}\in\mathbb{C}^{N_{s}\times N_{s}}$. Then, we can obtain the structure of the optimal solution to the problem (40) in the following Lemma.
@@
$\mathbf{\tilde{\Phi}}\triangleq{{\mathbf{U}}_{\text{B}}}\mathbf{\Sigma}_{\text{B}}^{-\frac{1}{2}}\mathbf{U}_{\text{B}}^{H}\left(\alpha_{1}^{2}{{\mathbf{\Phi}}_{1}}-{{\Gamma}_{s}}\sum\nolimits_{q=2}^{Q}{\alpha_{q}^{2}{{\mathbf{\Phi}}_{q}}}\right){{\mathbf{U}}_{\text{B}}}\mathbf{\Sigma}_{\text{B}}^{-\frac{1}{2}}\mathbf{U}_{\text{B}}^{H}.$ (45)
@@
$\nabla_{\mathbf{b}}f=-2\left(\left(\mathbf{I}+\mathbf{{\tilde{\Sigma}}}\mathbf{{{{\tilde{\Sigma}}}}^{H}}\right)^{-1}_{1:N_{s},1:N_{s}}\right)\mathbf{b}+\frac{2}{t}\left(\frac{\mathrm{diag}{{\left({{{\mathbf{\tilde{V}}}}^{H}}\mathbf{\tilde{B}\tilde{V}}\right)}_{1:{{N}_{s}}}}}{\left(\frac{{{N}_{s}}}{M}-\text{tr}\left({{{\mathbf{\tilde{V}}}}^{H}}\mathbf{\tilde{B}\tilde{V}\tilde{\Sigma}}{{{\mathbf{\tilde{\Sigma}}}}^{H}}\right)\right)}-\frac{\mathrm{diag}{{\left({{{\mathbf{\tilde{V}}}}^{H}}\mathbf{\tilde{\Phi}\tilde{V}}\right)}_{1:{{N}_{s}}}}}{\left(\text{tr}\left({{{\mathbf{\tilde{V}}}}^{H}}\mathbf{\tilde{\Phi}\tilde{V}\tilde{\Sigma}}{{{\mathbf{\tilde{\Sigma}}}}^{H}}\right)-{{\Gamma}_{0}}\right)}\right)\mathrm{diag}(\mathbf{b})$ (50)
@@
Please refer to [26, Prop. 7]. ∎
@@
The computational complexity of Algorithm 1 is dominated by the SVD decomposition of $\mathbf{B}\triangleq\tilde{\mathbf{U}}^{H}\mathbf{H}_{c}^{H}\mathbf{H}_{c}\tilde{\mathbf{U}}$, the calculation of Riemannian and Euclidean gradients, and the Riemannian projection in each iteration. Computing $\mathbf{B}$ has a complexity of $\mathcal{O}(N_{\text{RF}}N_{c}N)$, while the SVD of $\mathbf{B}$ requires $\mathcal{O}(N_{\text{RF}}^{3})$ operations [27]. The gradient calculations involve matrix multiplications and inversions with a total complexity of $\mathcal{O}(N_{\text{RF}}^{3}+N_{\text{RF}}^{2}N_{s}+N_{\text{RF}}{N_{s}}^{2})$. The Riemannian gradient descent update and projection require $\mathcal{O}(N_{\text{RF}}^{2})$ and $\mathcal{O}(N_{\text{RF}}^{3})$ operations, respectively [28]. Thus, the overall complexity of Algorithm 1 is $\mathcal{O}(I(N_{\text{RF}}N_{c}N+N_{\text{RF}}^{3}+N_{\text{RF}}^{2}N_{s}+N_{\text{RF}}{N_{s}}^{2}))$, where $I$ is the number of iterations.
Algorithm 1 Riemannian Joint Gradient Descent Algorithm
1:  Initialize $N_{s}$, $M$, ${\Gamma}_{s}$, $t>0$, the tolerances $\epsilon_{\tilde{\mathbf{V}}}>0$ and $\epsilon_{\mathbf{b}}>0$, the maximum number of iterations $I^{max}$, and choose a feasible $\mathbf{\tilde{V}}^{(0)}\in\mathbb{C}^{N_{\text{RF}}\times N_{\text{RF}}}$ such that $(\mathbf{\tilde{V}}^{(0)})^{H}\mathbf{\tilde{V}}^{(0)}=\mathbf{I}$ and $\mathbf{b}^{(0)}\in\mathbb{R}^{r}$.
@@
3:  while $n\leq I^{max}$ and $\|\xi_{\mathbf{\tilde{V}}}^{(n)}\|_{F}^{2}\geq\epsilon_{\tilde{\mathbf{V}}}$ or $\|\xi_{\mathbf{b}}^{(n)}\|^{2}\geq\epsilon_{\mathbf{b}}$ do

### v2

@@ III-C Problem Formulation
We jointly optimize the transmit and receive beamformers to maximize the achievable communication spectral efficiency while satisfying the following constraints: (i) the total transmit power budget, (ii) the hardware limitations imposed by the group-connected architecture and constant-modulus phase shifters, and (iii) the minimum required sensing SCNR for reliable target detection. Based on (2), (8) and (15), the optimization problem can be given by
$\max\limits_{{{\mathbf{W}}_{\text{BB}}},{{\mathbf{W}}_{\text{RF}}},\mathbf{w}}\!$ $\!\log\det\!\left(\!\mathbf{I}\!+\!\frac{1}{\sigma_{c}^{2}}{{\mathbf{H}}_{c}}{{\mathbf{W}}_{\text{RF}}}\!{{\mathbf{W}}_{\text{BB}}}\!\mathbf{W}_{\text{BB}}^{H}\mathbf{W}_{\text{RF}}^{H}\mathbf{H}_{c}^{H}\!\!\right)$ (17a)
@@
$\frac{{{\mathbf{w}}^{H}}{{\mathbf{G}}_{0}}\left({{\mathbf{W}}_{\text{RF}}}{{\mathbf{W}}_{\text{BB}}}\mathbf{W}_{\text{BB}}^{H}\mathbf{W}_{\text{RF}}^{H}\right)\mathbf{G}_{0}^{H}\mathbf{w}}{{{\mathbf{w}}^{H}}\left(\sum\nolimits_{q=1}^{Q-1}{{{\mathbf{G}}_{q}}\left({{\mathbf{W}}_{\text{RF}}}{{\mathbf{W}}_{\text{BB}}}\mathbf{W}_{\text{BB}}^{H}\mathbf{W}_{\text{RF}}^{H}\right)\mathbf{G}_{q}^{H}}+\sigma_{s}^{2}{{\mathbf{I}}_{N}}\right)\mathbf{w}}\geq{{\Gamma}_{s}},$ (17d)
where ${{\mathcal{A}}_{F}}$ is the set of block matrices, with each block being an $M\times M_{\text{RF}}$ matrix with constant-magnitude entries. However, due to the specific block-diagonal structure of the analog beamformer in ${\mathcal{A}}_{F}$, this type of joint analog-digital optimization problem is non-convex and intractable.
@@ IV Low-Complexity Hybrid Beamforming Design
In this section, we first analyze the optimal transmit covariance matrix structure and obtain the closed-form analog beamformer. Building upon this, we propose two efficient algorithms to optimize the digital beamformer: a joint Riemannian-Euclidean gradient descent method and an SDR-based randomization approach.
@@ IV-A The Optimal Waveform Covariance Matrix
To provide an optimal solutions to the problem (18), we propose to exploit the structure of the communication channel and sensing array response vectors based on the piecewise-far-field channel model, respectively. Namely, we leverage the following observations:
IV-A1 Structure of the communication channel
@@
IV-A2 Structure of Sensing Array Response Vectors
The subarray response vectors $\mathbf{a}_{tq}^{1}(\mathbf{l}q),\ldots,\mathbf{a}_{tq}^{K}(\mathbf{l}_{q})$ for a given sensing object $q$ are linearly independent due to the widely-spaced subarray configuration, which results in different subarrays observing the target from distinct AoD. Therefore, each sensing array response vector ${{\mathbf{g}}_{tq}}$ can be expressed as a fixed linear combination of $K$ linearly independent vectors as
$\mathbf{\bar{a}}_{tq}^{k}{{\left(\phi_{tq}^{k}\right)}}\triangleq[\ \underbrace{\mathbf{0},\ldots,\mathbf{0}}_{k-1},\mathbf{a}_{tq}^{k}{{\left(\phi_{tq}^{k}\right)}^{H}},\underbrace{\mathbf{0},\ldots,\mathbf{0}}_{K-k}\ ],\in\mathbb{C}^{1\times KM},$ (20)
where $k=1,2,\cdots,K$, $q=0,1,\cdots,Q-1$, and $\mathbf{0}$ is an all-zero vector of dimension $1\times M$.
@@ IV-A2 Structure of Sensing Array Response Vectors
${{\mathbf{g}}_{tq}}={{\mathbf{\bar{A}}}_{tq}}{\bm{\nu}}_{tq},\ \forall q,$ (21)
@@ IV-A3 Joint Representation of Communication Channel and Sensing Array Response Vectors
${{\mathbf{\tilde{U}}}_{k}}\!$ $=\!\left[\mathbf{\bar{a}}_{t0}^{k}\!\left(\phi_{t0}^{k}\right)\!,\ldots,\mathbf{\bar{a}}_{tQ-1}^{k}\!\left(\!\phi_{tQ-1}^{k}\!\right)\!,\mathbf{\tilde{a}}_{t1}^{k}\!\left(\theta_{t1}^{k}\right)\!,\ldots,\mathbf{\tilde{a}}_{t{{N}_{p}}}^{k}\!\left(\!\theta_{t{{N}_{p}}}^{k}\!\right)\!\right],$ (30a)
${{\mathbf{A}}_{kk}}\!$ $=\!\left[\mathbf{a}_{t0}^{k}\!\left(\phi_{t0}^{1}\right)\!,\ldots,\mathbf{a}_{tQ-1}^{k}\!\left(\phi_{tQ-1}^{1}\right)\!,\mathbf{a}_{t1}^{k}\!\left(\theta_{t1}^{k}\right)\!,\ldots,\mathbf{a}_{t{{N}_{p}}}^{k}\!\left(\!\theta_{t{{N}_{p}}}^{k}\!\right)\!\right].$ (30b)
@@ IV-B Equivalent Low-dimensional Optimization Problem
To fully exploit the spatial multiplexing gain, we transmit $N_{s}=\mathrm{rank}(\mathbf{H}_{c})$ independent data streams to the user. In addition, in the mmWave/sub-THz band, where reflection and scattering losses are significant, the contributions of high-order reflection and scattering paths can be neglected, leading to a small value of $N_{p}$. Therefore, in scenarios where the number of subarrays and the number of interferences are limited, it is reasonable to consider that the available RF chains are sufficient and exceed $K(N_{p}+Q)$. Consequently, to enable sufficient spatial DoFs for both communication and sensing, we activate $M_{\text{RF}}=(N_{p}+Q)$ RF chains per subarray, yielding a total of $N_{\text{RF}}=K(N_{p}+Q)$ RF chains in the system.
@@
${{\gamma}_{s}}=\frac{\alpha_{1}^{2}\text{tr}\left({{\mathbf{W}}_{\text{BB}}}\mathbf{W}_{\text{BB}}^{H}{{\mathbf{\Phi}}_{0}}\right)}{\sum\nolimits_{q=1}^{Q}{\alpha_{q}^{2}\text{tr}\left({{\mathbf{W}}_{\text{BB}}}\mathbf{W}_{\text{BB}}^{H}{{\mathbf{\Phi}}_{q}}\right)}+\sigma_{s}^{2}\mathbf{w}{{\mathbf{w}}^{H}}},$ (33)
where ${{\mathbf{\Phi}}_{q}}\triangleq{{\mathbf{\tilde{U}}}^{H}}{{\mathbf{g}}_{tq}}\mathbf{g}_{rq}^{H}\mathbf{w}{{\mathbf{w}}^{H}}{{\mathbf{g}}_{rq}}\mathbf{g}_{tq}^{H}\mathbf{\tilde{U}},q\in\left\{0,\ldots,Q-1\right\}$. Moreover, due to the block diagonal structure of $\mathbf{{W}}_{\text{RF}}$, each non-zero element of $\mathbf{W}_{\text{RF}}$ is multiplied with the corresponding row in $\mathbf{{W}}_{\text{BB}}$. Hence, the constraint (17b) can be simplified as
@@
$\text{s}.\text{t}.\ \ \text{tr}\left({{\mathbf{W}}_{\text{BB}}}\mathbf{W}_{\text{BB}}^{H}\right)\leq\frac{{{N}_{s}}}{M},$ (35b)
$\ \ \ \alpha_{0}^{2}\text{tr}\left(\mathbf{W}_{\text{BB}}^{H}{{\mathbf{\Phi}}_{0}}{{\mathbf{W}}_{\text{BB}}}\right)\text{-}{{\Gamma}_{s}}\sum\limits_{q=1}^{Q-1}{\alpha_{q}^{2}\text{tr}\left(\mathbf{W}_{\text{BB}}^{H}{{\mathbf{\Phi}}_{q}}{{\mathbf{W}}_{\text{BB}}}\right)}\geq{{\Gamma}_{0}},$ (35c)
where $\Gamma_{0}\triangleq{{\Gamma}_{s}}\sigma_{s}^{2}\mathbf{w}{{\mathbf{w}}^{H}}$. Note that the constraint (35c) is an equivalent transformation of the sensing SCNR constraint in (17d) based on (33).
Although the digital beamformer optimization problem (35), has a reduced dimension, it remains non-convex due to the rank constraint imposed by the limited number of data streams. To tackle this non-convex problem, we propose two distinct algorithms in the following subsections: a manifold optimization method that directly optimizes the digital beamformer on the rank-constrained space and an SDR-based method that obtains a near-optimal solution. By investigating these two optimization strategies, we provide a comprehensive framework for solving the rank-constrained digital beamformer design problem.
@@ IV-C Joint Optimization on Riemannian Manifold and Euclidean Space
In this subsection, we first derive the structure of the optimal solution to problem (35) without relaxing the rank constraint, and obtain the semi-closed-form solution. Based on this, we then develop a Riemannian joint gradient descent algorithm.
Let $\mathbf{B}\triangleq{{\mathbf{\tilde{U}}}^{H}}\mathbf{H}_{c}^{H}{{\mathbf{H}}_{c}}\mathbf{\tilde{U}}$, and it follows that $\mathbf{B}$ is a Hermitian matrix with $\mathrm{rank}(\mathbf{B})=\min\{N_{\text{RF}},N_{s}\}=N_{s}$. Performing eigendecomposition on $\mathbf{B}$ and retaining only the non-zero eigenvalues and their corresponding eigenvectors, we obtain $\mathbf{B}={{\mathbf{U}}_{B}}{{\mathbf{\Sigma}}_{B}}\mathbf{U}_{B}^{H}$, where ${{\mathbf{U}}_{B}}\in\mathbb{C}^{N_{\text{RF}}\times N_{s}}$, and $\mathbf{\Sigma}_{B}\in\mathbb{C}^{N_{s}\times N_{s}}$. Then, we can obtain the structure of the optimal solution to the problem (35) in the following Lemma.
@@
$\mathbf{\tilde{\Phi}}\triangleq{{\mathbf{U}}_{\text{B}}}\mathbf{\Sigma}_{\text{B}}^{-\frac{1}{2}}\mathbf{U}_{\text{B}}^{H}\left(\alpha_{0}^{2}{{\mathbf{\Phi}}_{0}}-{{\Gamma}_{s}}\sum\nolimits_{q=1}^{Q-1}{\alpha_{q}^{2}{{\mathbf{\Phi}}_{q}}}\right){{\mathbf{U}}_{\text{B}}}\mathbf{\Sigma}_{\text{B}}^{-\frac{1}{2}}\mathbf{U}_{\text{B}}^{H}.$ (40)
@@
$\nabla_{\mathbf{b}}f=-2\left(\left(\mathbf{I}+\mathbf{{\tilde{\Sigma}}}\mathbf{{{{\tilde{\Sigma}}}}^{H}}\right)^{-1}_{1:N_{s},1:N_{s}}\right)\mathbf{b}+\frac{2}{t}\left(\frac{\mathrm{diag}{{\left({{{\mathbf{\tilde{V}}}}^{H}}\mathbf{\tilde{B}\tilde{V}}\right)}_{1:{{N}_{s}}}}}{\left(\frac{{{N}_{s}}}{M}-\text{tr}\left({{{\mathbf{\tilde{V}}}}^{H}}\mathbf{\tilde{B}\tilde{V}\tilde{\Sigma}}{{{\mathbf{\tilde{\Sigma}}}}^{H}}\right)\right)}-\frac{\mathrm{diag}{{\left({{{\mathbf{\tilde{V}}}}^{H}}\mathbf{\tilde{\Phi}\tilde{V}}\right)}_{1:{{N}_{s}}}}}{\left(\text{tr}\left({{{\mathbf{\tilde{V}}}}^{H}}\mathbf{\tilde{\Phi}\tilde{V}\tilde{\Sigma}}{{{\mathbf{\tilde{\Sigma}}}}^{H}}\right)-{{\Gamma}_{0}}\right)}\right)\mathrm{diag}(\mathbf{b}).$ (45)
@@
Please refer to [28, Proposition 7]. ∎
@@
The computational complexity of Algorithm 1, hereafter referred to as Riemannian-Euclidean Joint Gradient Descent (RM-JGD), is dominated by the SVD decomposition of $\mathbf{B}\triangleq\tilde{\mathbf{U}}^{H}\mathbf{H}_{c}^{H}\mathbf{H}_{c}\tilde{\mathbf{U}}$, the calculation of Riemannian and Euclidean gradients, and the Riemannian projection in each iteration. Computing $\mathbf{B}$ has a complexity of $\mathcal{O}(N_{\text{RF}}N_{c}N)$, while the SVD of $\mathbf{B}$ requires $\mathcal{O}(N_{\text{RF}}^{3})$ operations [29]. The gradient calculations involve matrix multiplications and inversions with a total complexity of $\mathcal{O}(N_{\text{RF}}^{3}+N_{\text{RF}}^{2}N_{s}+N_{\text{RF}}{N_{s}}^{2})$. The Riemannian gradient descent update and projection require $\mathcal{O}(N_{\text{RF}}^{2})$ and $\mathcal{O}(N_{\text{RF}}^{3})$ operations, respectively [30]. Thus, the overall complexity of Algorithm 1 is $\mathcal{O}(I(N_{\text{RF}}N_{c}N+N_{\text{RF}}^{3}+N_{\text{RF}}^{2}N_{s}+N_{\text{RF}}{N_{s}}^{2}))$, where $I$ is the number of iterations.
Algorithm 1 RM-JGD Algorithm
1:  Initialize $N_{s}$, $M$, ${\Gamma}_{s}$, $t>0$, the tolerances $\epsilon_{\tilde{\mathbf{V}}}>0$ and $\epsilon_{\mathbf{b}}>0$, the maximum number of iterations $I_{\max}$, and choose a feasible $\mathbf{\tilde{V}}^{(0)}\in\mathbb{C}^{N_{\text{RF}}\times N_{\text{RF}}}$ such that $(\mathbf{\tilde{V}}^{(0)})^{H}\mathbf{\tilde{V}}^{(0)}=\mathbf{I}$ and $\mathbf{b}^{(0)}\in\mathbb{R}^{r}$.
@@
3:  while $n\leq I_{\max}$ and $\|\xi_{\mathbf{\tilde{V}}}^{(n)}\|_{F}^{2}\geq\epsilon_{\tilde{\mathbf{V}}}$ or $\|\xi_{\mathbf{b}}^{(n)}\|^{2}\geq\epsilon_{\mathbf{b}}$ do